OpenStudy (anonymous):

Solve the system:2x – 5y = 1 3x – 4y = -2 write your answer as an ordered pair, (x,y).

OpenStudy (anonymous):

i got y=1 i got confused after that with finding x can someone help me

OpenStudy (anonymous):

The answer is (-2,-1) Substitute the value of y to the 1st equation 2x-5y=1 2x-5(-1)=1 2x=1-5 2x=-4 X=-2

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Your welcome :)